Solve system of equations
X+3y-2z=-9
6x-4y+3z=29
2x+2y+z=13

Answer by ewatrrr(24785)


Hi
I. x+3y-2z=-9
II. 6x-4y+3z=29
III. 2x+2y+z=13
solving by elimination
I and III
x+3y-2z=-9
4x+4y+2z=26
5x + 7y = 17
II and III
6x-4y+3z=29
-6x-6y-3z=-39
-10y = -10
y = 1 and x = 2 | 5x = 17-7y, x = 10/5
x+3y-2z=-9
OR x+3y+9 = 2z
14 = 2z z = 7